Solve the following recurrence relations
i) Fn= Fn-1 +Fn-2 where a1=a2=1
ii) an=2an-1 - an-2 +2 where a1 = 1 and a2 = 5
1
Expert's answer
2020-11-30T10:37:25-0500

i) Write down the characteristic equation:

"x^2 = x + 1"

"x^2 -x -1 =0"

"x_1 =\frac{1 + \sqrt5}2 = \phi ; \ x_2 = \frac {1 -\sqrt 5}2 =\psi"

"F_n = ax_1^n +bx_2^n"

"a*\phi +b*\psi = 1\ and\ \ a*\phi^2 +b*\psi^2 = 1"

"a*\phi = 1 - b*\psi;"

"(1-b*\psi)*\phi +b*\psi^2 = 1"

"\phi*\psi = \frac{1- 5}4 = -1;"

"\phi +b + b*\psi^2 = 1"

"b*(1+\psi^2) = 1 -\phi"

"\psi^2 > 0;\ \ 1+ \psi^2 >0"

"b = \frac {1-\phi}{1+\psi^2} = \frac {1- \frac{1+\sqrt 5}2}{1+ (\frac{1-\sqrt5}2)^2} ="

"= \frac{\frac{1-\sqrt5}{2}}{\frac{2*5 -2*\sqrt5}4} = \frac{\frac{1-\sqrt5}{2}}{-\sqrt5*\frac{(1 - \sqrt5)}2}=-\frac {1}{\sqrt5}"

"\phi > 0;\ \ a = \frac{1 + \frac{\psi}{\sqrt5}}{\phi} =\frac{1+\frac{\frac{1-\sqrt5}{2}}{\sqrt5}}{\frac{1+\sqrt5}{2}}="

"=\frac{\frac{2*\sqrt5+1-\sqrt5}{2\sqrt5}}{\frac{1+\sqrt5}{2}} =\frac{\frac{1+\sqrt5}{2\sqrt5}}{\frac{1+\sqrt5}{2}}="

"=\frac{1}{\sqrt5}*\frac{\frac{1+\sqrt5}{2}}{\frac{1+\sqrt5}{2}} = \frac{1}{\sqrt5}"

"F_n = \frac{\phi^n}{\sqrt5} - \frac{\psi^n}{\sqrt5}= \frac{\phi^n - \psi^n}{\sqrt5};"


ii) "a_3 = 2*5 -1 + 2 = 10 + 1 = 11"

"a_{n+1} = 2a_{n} - a_{n-1} + 2;"

"a_n = 2a_{n-1} - a_{n-2} + 2;"

From the first equation subtract the second:

"a_{n+1} - a_n = 2a_n - a_{n-1} + 2 - 2a_{n-1} + a_{n-2} - 2"

"a_{n+1} = 3a_n - 3a_{n-1} + a_{n-2} \ for \ n \ge 3"

Therefore, for "n \ge 4: a_n = 3a_{n-1} - 3a_{n-2} + a_{n-3};"

Write down the characteristic equation:

"x^3 - 3x^2 +3x -1 = 0"

"(x-1)^3= 0"

"x_{1,2,3}= 1;"

"a_n = k_1*1^n + k_2*n*1^n + k_3*n^2*1^n ="

"=k_1+ k_2*n + k_3*n^2;"


The system of three equations:

"k_1+ k_2 + k_3= a_1 = 1"

"k_1+ k_2*2 + k_3*4= a_2 = 5"

"k_1+ k_2*3 + k_3*9= a_3 = 11"


Subtract the first from the second:

"k_2 + k_3*3= 4" (iv)


Subtract the second from the third:

"k_2 + k_3*5= 6" (v)


Subtract (iv) from (v):

"k_3*2 = 2"

"k_3 = 1"


From this and (iv):

"k_2 +3 = 4"

"k_2 = 1"


From these and the first:

"k_1 + 1 +1 = 1"

"k_1=-1"


"a_n = -1 + n + n^2 = n^2 +n -1."


Answer: i) "F_n = \frac{\phi^n - \psi^n}{\sqrt5}"

ii) "a_n = n^2 +n -1"